Circular No. 5626

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SLAUGHTER-BURNHAM (1992w) This comet has been recovered by S. M. Larson and C. D. Slaughter

with the 1.5-m reflector (+ CCD) at the Lunar and Planetary Laboratory’s Catalina Station. Co-adding of the Cousins R-band images showed a 10′′ coma and hint of a tail. The comet was also detected on Sept. 18. The following positions are very close to the prediction on MPC 16380 (ephemeris on MPC 19888).

1992 UT  α_{2000}  δ_{2000}  m_{1}
Sept.24.18228  20^{h}46^{m}17.^{s}20  −22^{o}22′52′′.2  21.5
24.18751  20 46 17.19  −22 22 50.6
24.19206  20 46 17.18  −22 22 51.0
24.20947  20 46 16.94  −22 22 48.7
25.13324  20 46 09.48  −22 20 30.2
25.13939  20 46 09.44  −22 20 29.0
25.15280  20 46 09.31  −22 20 27.6
25.15962  20 46 09.28  −22 20 26.2
25.16830  20 46 09.16  −22 20 24.4
25.17709  20 46 09.11  −22 20 23.2

SUPERNOVA 1992au IN ANONYMOUS GALAXY M. M. Phillips, Cerro Tololo Interamerican Observatory, reports that

a spectrogram of SN 1992au obtained with the CTIO 4.0-m telescope on Oct. 1.17 UT shows that this object is a type-Ia supernova. Comparison with spectra of the prototype-Ia SN 1989B indicates that maximum light for SN 1992au probably occurred 70 ± 10 days ago, consistent with this supernova having been discovered near peak brightness (see IAUC 5584). A spectrum obtained of the host galaxy of SN 1992au yields a redshift of 0.062.

NOVA SCORPII 1992 Additional photometry (cf. IAUC 5574, 5584) by A. C. Gilmore, Mt.

John University Observatory (uncertainty ± 0.02): Sept. 17.39 UT, V = 10.47, B–V = +0.81, U–B = +0.02, V–R = +0.96, V–I = +1.49.

Further unfiltered CCD magnitudes by W. Liller, Viña del Mar, Chile: Sept. 2.01 UT, 8.76; 2.97, 8.99; 5.00, 9.12; 6.01, 9.33; 7.00, 9.44 (comparison star SAO 227635).

1992 October 2 Daniel W. E. Green